Service Observing for DTime UT Date 06072016

On site: David, Barry

Tucson: Jenny, Olga, Shane, Andrew, John H., Michelle DLP, Doug, Kellee.



Local time (MST) used throughout.

19:31 Sunset

20:10 David is correcting pointing and collimating.

20:23 Going to telluric for Barry\x92s first target.

20:28 12 deg twilight.

20:29 0\x94.9 - 1\x94 on guider.

20:31 DX LUCI WFS looks a little odd. John suspects a bad flat.

20:36 Using LS_2.0_1.5, centering in 1\x94.5 slit. (Lower slit as displayed.) Centered target on 1009.24, 517.01. Offset required was +7\x94.51, -59\x94.45.

20:40 Starting telluric observation.

20:43 Going to science target.

20:45 Starting acquisition.

20:48 LUCI2 error \x91no mask used\x92...

20:52 Having double checked the script we are calling Dave T.

21:05 Michelle leaves for the night.

21:10 Making some very small rotator changes and LUCI2 detector reads while Dave works. (To test the acquisition rotation sign issue.)

21:18 Looks like LUCI2 is out of action for the moment. Switching to LBCs.

21:50 moved to LBC first target: Bootes1, transiitting

22:52 Ran test of wrs but got error.

% WRITE_JPEG: Error opening file. Unit: 102

File: save/FINDPUP_results_2016Jun 7_04:50:15.jpg

No such file or directory

23:00 Running dofpia on Bootes1

23:10 dofpia hung, the data files never appeared in newdata.

23:16 Running dofpia again. Z22 runaway

23:37 Running dofpia again. Z22 on LBCB continues to runaway. Quit it and allowed observations to proceed

23:47 Bootes 2 started, LBCB ~ 1.3\x94, LBCR ~0.8\x94

22:35 Struggling with dofpia. It thinks the seeing is too bad on the blue side to correct the spherical, but red has worked fine.

22:44 Running the co-pointing script.

22:53 Going to Bootes 1.

23:09 Dimm is ~1\x94, LBCB images showing around 1.4\x94 in USpec , but LBCR ~1.0\x94 in RBessel

23:18 Running dofpia again

00:16 Switching to Christians targets. Large gradient on SX, at 44 degrees elevation and sinking.

0:22 Running copointing and dofpia on Veillet OB1

0:40 Veillet OB2 Slew & Execution, LBCR ~1.2\x94, LBCB ~1.4\x94

0:55 Veillet OB3 Slew & Execution, LBCR ~1.0\x94, LBCB ~1.3\x94

01:06 Veillet OB4 Slew & Execution, LBCR ~1.1\x94, LBCB ~1.4\x94

01:11 Veillet OB5 Slew & Execution, LBCR ~ 1.2\x94, LBCB ~ 1.5\x94

01:16 Veillet OB6 Slew & Execution, LBCR ~ 1.2\x94, LBCB ~ 1.2 \x94

01:20 Reconfiguring to MODS1 MODS2

01:40 Problems with SX shell, David restarting

01:51 Back in business, pointing checks, collimating

02:19 Slew to FF1725+3709, Aquisition

02:46 MODS2 FF1725+3709 science obs started, pointing adjustment made to MODS1 & check

02:49 MODS1 FF1725+3709 science obs started.

03:54 Slewing to standard, BD284211

03:55 SX primary panic

04:02 Active optics corrections removed to help with mirror recovery.

04:08 12 deg twilight.

04:14 Starting science

04:25 Finish obs, closing. Powering off LBCs

04:41 Starting MODS1 & 2 calibrations: Grating Arcs, Slitless Flats, & Biases

05:06 Sunrise


05:28 Calibrations finished

Comments


 
Topic revision: r1 - 07 Jun 2016, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback